Commander Report – November 29th

Sol: 2
Summary Title: Experiment Preparation and EVAs 1&2
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 enjoyed another day of beautiful weather on Mars, with light high-altitude clouds in the morning, becoming clear and sunny in the afternoon. The crew enjoyed oatmeal, homemade sourdough bread, and espresso for breakfast. Technology Officer completed preparations for the afternoon EVA while the rest of the crew completed final preparations, training, and exploration of the local area. After a quick lunch, CDR, XO, and Green Hab Officer embarked on EVA #1 to Marble Ritual. The 1 hour EVA was nominal and conducted rover, analog spacesuit, and communications checkout. Upon the return of the EVA #1 crew, HSO, Technology Officer, and Engineer departed on EVA #2 to Marble Ritual, completed their own rover, analog spacesuit, and communications checkout, then continued on to the location across Cow Dung Rd. where the mirror samples were deployed yesterday. There the crew used a remote control mini-ATV with cameras, controlled from a follow distance of roughly 80-100 ft. The trio drove the mini-ATV from Marble Ritual, south along Cow Dung Rd., and back to the road which turns west back towards the crew habitat collecting data along the way. Upon return from EVA #2, the crew experienced an anomaly with the front outer airlock door and a minor anomaly with the plug-in attachments for one of the space suits. Details of these anomalies are covered in the Operations Report and EVA Reports from today. In the evening, the crew set up the Aerospace microscope in the Science Dome and finished out Green Hab training and the final sections of the HSO checklist. The crew enjoyed spaghetti with meat sauce and homemade garlic bread for dinner and discussed their personal goals for the mission and agreed upon their specific simulation rules.

Copyright © The Mars Society. All rights reserved. | Main Site